# Statistical Validation of Contagion Centrality in Financial Networks

### v1

@@ Abstract
In this paper, we introduce a novel centrality measure to evaluate shock propagation on financial networks capturing a notion of contagion and systemic risk contributions. In comparison to many popular centrality metrics (e.g., eigenvector centrality) which provide only a relative centrality between nodes, our proposed measure is in an absolute scale permitting comparisons of contagion risk over time. In addition, we provide a statistical validation method when the network is estimated from data, as is done in practice. This statistical test allows us to reliably assess the computed centrality values. We validate our methodology on simulated data and conduct empirical case studies using financial data. We find that our proposed centrality measure increases significantly during times of financial distress and is able to provide insights in to the (market implied) risk-levels of different firms and sectors.
Keywords: Finance, Contagion Risk, Financial Stability, Network Analysis, Statistical Validation
@@ 1 Introduction
The growing complexity of financial systems requires the creation of a framework that can assess the interconnections among firms, see e.g. (Feinstein and Hałaj, , 2023, Figure 1). These connections play a crucial role in capturing systemic events and shocks that can lead to system instability. Researchers have shown growing interest in modeling financial system using network representations to gain a deeper understanding of these effects (Huang et al., , 2013; Jackson and Pernoud, , 2020). Emerging as a significant driver of instability, contagion, often described as the propagation of information or shocks across nodes, results in cascading effects that spread throughout the entire network, akin to an epidemic (Jalili and Perc, , 2017). Within this paper, we construct a novel network centrality measure designed to capture financial contagion. Due to the proposed construction, we are able to, e.g., statistically test our centrality measure; as far as the authors are aware, this is the first network centrality that can be tested.
The structure of this paper is as follows: In this section, we provide an introduction to the research topic with relevant literature. Section 2 outlines the approach used to construct the network and introduces the Leontief inverse centrality measure, referred to as the “Leontief Centrality Measure (LCM)” hereafter. With this definition, we detail the statistical framework for the LCM in Section 3 and validate this framework using simulated data. We apply the LCM approach to financial data scenarios in Section 4 to assess framework’s applicability to capture stress events. Finally, in Section 5, we summarize the key findings of the paper and propose potential avenues for future research.
1.1 Literature Review
@@
Due to the recent financial events, it is important to view the financial system as a complex network, where firms serve as nodes and financial dependencies as links (Battiston et al., , 2012). There are multiple perspectives on financial networks, e.g., (i) informational (Fink et al., , 2016), (ii) interbank or default contagion (Acemoglu et al., , 2016), (iii) portfolio overlap or price-mediated contagion (Cont and Schaanning, , 2017), and (iv) cross-ownership (Fichtner et al., , 2017). The construction of financial networks involves various methods, including correlation (Pacreau et al., , 2021), partial correlation (Millington and Niranjan, , 2020), and analysis of financial dependencies based on balance sheets and contracts between institutions (Eisenberg and Noe, , 2001).
@@
Failures or contagion can be triggered by common factors, such as market declines, or by idiosyncratic shocks affecting individual institutions, like spread widening (Benoit et al., , 2017). The consequences of a shock can be broken down into two parts: the immediate failure of an institution triggered by a sufficiently large shock, and the spread of the shock throughout the system due to network effects. While many studies focus on the direct collapse of a firm (fundamental default) resulting from an external shock, they often overlook the aggregated impact of network-induced contagion (default by contagion) as noted by Cont et al., (2013). It is important to distinguish between the first-order effect, which could be the default of a firm or the occurrence of a shock impacting the firm, and the higher-order impacts due to these stresses. As highlighted by di Iasio et al., (2013); Veraart, (2020), distress contagion can propagate through the network, causing widespread losses, even when there are no defaults within financial system. To measure these effects, it is essential to define a suitable centrality measure.
@@
Centrality measures serve as valuable tools to quantitatively assess the structural significance of nodes within a network. As stated by D’Arcangelis and Rotundo, (2016), a higher centrality score indicates that a particular property is more fitting for a given node. Different centrality measures offer distinct insights into various centrality dimensions (Hevey, , 2018). In addition, the centrality notion becomes more crucial when the networks are not symmetric (Capponi and Larsson, , 2015) which is the case in financial systems. When examining propagation within networks, Pacreau et al., (2021) highlights that the centrality of a node indicates its sensitivity to market fluctuations and reveals its connectivity within the network. Nodes with high centrality are often connected to many other nodes or are neighbors to densely interconnected subnetworks. Along this line, Battiston et al., (2012) introduces DebtRank, utilizing feedback-centrality to measure the impact of a shock on a single node. Building on this concept, Puliga et al., (2014) presents Group Debtrank to quantify the centrality of nodes when small shocks affect firms and concludes that during crisis periods, the centrality measure tends to increase.
@@
It is crucial to select an appropriate centrality measure that aligns with the attribute under consideration (Singh et al., , 2020). In this work, we introduce a simple yet innovative centrality measure that reflects the spread of a shock over the network structure based on the Leontief inverse (Leontief, , 1936, 1986). Widely used in input-output analysis in economics (Antràs et al., , 2012; Antràs and Chor, , 2013, 2018), the Leontief inverse has also found applications in network science more recently (Moran and Bouchaud, , 2019). Certain studies in the financial network field have employed the Leontief inverse to establish the network structure directly (Elliott et al., , 2014; D’Arcangelis and Rotundo, , 2016), rather than applying it to compute the centrality.
@@ 1.2 Primary Contributions
In this paper, we construct weighted directed networks using a vector autoregressive (VAR) model, where nodes represent firms and the links depict interconnections between them. Rather than utilizing the Leontief inverse notion solely as a generative method for networks, we redefine it as the Leontief centrality measure, broadening its applicability across different network structures. This measure, referred to as LCM, quantifies node centrality during shock scenarios, illustrating a node’s position when hit by a shock. It considers the propagation and contagion risks inherent in the network’s topology, and boils down the shock propagation dynamics to a closed form.
We define LCM at pair-, node-, and system-levels, offering different granularities for analysis. Pair-level LCM between two firms demonstrates the first firm’s position when the second firm experiences a shock. Node-level LCM of a firm is the total resulting shock when it is stressed. System-level LCM aggregates multiple scenarios to portray overall centrality changes under equally probable situations. We also allow the association of weights with node- and system-level centralities.
Our approach involves statistical analysis to assess the robustness of the centrality measure. We derive the asymptotic distribution of LCM through its Taylor expansion and form hypotheses for testing – specifically focusing on whether LCM significantly differs from zero. We validate our framework using simulated data, demonstrating the theoretical distribution’s compatibility with empirical observations. In this simulated setup, the true value of LCM consistently falls within the confidence intervals of the theoretical distribution at a 95% confidence level.
Applicability in financial contexts is a key aspect of our paper. We evaluate the performance of our framework using two financial datasets – Credit Default Swap Index (CDX) and equity tick data. In the CDX case study, LCM effectively captures underlying market dynamics, outperforming both degree centrality and the leading eigenvalue. In the case of tick data, LCM provides insights into the market and institutions’ states but does not suffice for standalone investment decisions. However, a blend of node- and system-level LCM dynamics along with institutional background information can guide investment strategies effectively.
@@ 2 Network Centrality
A network consists of a set of nodes and links. The connectivity pattern between nodes is captured by the adjacency matrix $A\in\mathbb{R}^{M\times M}$, where $M$ denotes the number of nodes, and its elements $a_{ij}$, for $i,j=1,...,M$, indicate the presence or absence of connections between nodes. For our purpose, the elements $a_{ij}$ represent the connectivity weights between nodes $i$ and $j$ (Papana et al., , 2017).
In this paper, we consider the nodes to be the firms and the links to be the connections based on a linear model. Our network construction follows a VAR($l$) structure considering only lag $l$. The structure is as follows:
$\begin{gathered}X_{t}=X_{t-l}A+\mathcal{E}_{t}\,,\end{gathered}$ (1)
where $X_{t}\in\mathbb{R}^{M}$ is a vector (of, e.g., market returns of each firm) at time $t$, while $\mathcal{E}_{t}\in\mathbb{R}^{M}$ denotes the independent and identically distributed (iid) noise. As a weighted directed matrix, $A$ serves as the adjacency matrix of the network, showcasing the interconnectedness of different variables. We refer to this system as the linear model.
Assumption 1 (Stationarity).
@@
For the remainder of this work, we assume that the spectral norm of the adjacency matrix is strictly less than one, i.e., the VAR($l$) model is stationary.
Remark 1.
If one wants to explore the idiosyncratic relations among time series, dimensionality reduction to eliminate the common factors can be performed. The dimension reduction can be achieved using various methods such as Singular Value Decomposition (SVD) or spectral reduction (Ricciardi et al., , 2022).
@@
We introduce a novel centrality measure that overcomes the limitations of relative measures such as the eigenvector centrality (which can be scaled arbitrarily; see, e.g., Markose et al., (2012); Chan-Lau, (2018); Olmo, (2021)) to allow for direct comparison between different time windows. This measure defined as in Equation 2 below, is readily interpretable in the context of financial contagion.
As an example, in Figure 1 we present a schematic network diagram along with its shock propagation dynamics. Consider four time series, corresponding to four nodes denoted as $W,X,Y$ and $Z$, which are connected as illustrated in the figure on the left. The shock propagation pattern is shown in the figure on the right side, under the scenario in which node $W$ experiences shock $\epsilon$ at time $t=0$. Due to the connections with nodes $X$ and $Y$, this shock subsequently propagates and affects these nodes at time $t=1$. Following the network, the shock then cycles back to node $W$ and hits $Z$ at $t=2$. This cascading effect goes on until the shock dissipates.
[…]
In order to formulate these dynamics more generally, we assume $M$ time series and define an initial shock vector $\bm{\epsilon}\in\mathbb{R}^{M}$, where element $i$ represents the magnitude of the shock to node $i$. The shock propagation dynamics can be shown as:
$\begin{gathered}S_{t}=S_{t-1}A,\quad S_{0}=\bm{\epsilon}\quad\Rightarrow\quad S_{t}=\bm{\epsilon}A^{t},\end{gathered}$
where $S_{t}$ is the state of the system at time $t$ and $A$ is the adjacency matrix. By applying the adjacency matrix $A$ to each state iteratively, we can trace the propagation of the shock over time. Specifically, $S_{1}=S_{0}A$, $S_{2}=S_{1}A=S_{0}A^{2}$, and so on. Each term in the series $\bm{\epsilon}A^{t}$ represents the effect of the shock propagating through the network over different time intervals, accounting for higher-order interactions and contagion effects. The summation
$\sum_{t=1}^{\infty}S_{t}=\bm{\epsilon}A+\bm{\epsilon}A^{2}+...=\bm{\epsilon}\,\underbrace{(A+A^{2}+...)}_{\text{pair-level LCM}}$
captures the cumulative effect of the shock propagation and represents the centrality of the system. Each element in this matrix represents the pair-level LCM, i.e., the centrality of node $j$ (corresponding to the column index) when node $i$ (corresponding to the row index) is hit by a shock. Note that this pair-level LCM is independent of the initial shock $\bm{\epsilon}$.
@@
Under Assumption 1, the pair-level LCM is well-defined and can be simplified via the Leontief inverse:
$\sum_{t=1}^{\infty}A^{t}=(I-A)^{-1}-I\,.$ (2)
@@
LCM is not solely influenced by the strength of the links but also by the overall structure of the network. A densely connected network with comparatively weaker links may yield a higher centrality measure than a sparsely connected network with stronger links.
To compute the node-level centrality when node $i$ is shocked, we consider the weighted summation of each row of the centrality measure matrix:
$C=[\,(I-A)^{-1}-I\,]\,\bm{w}\,,$ (3)
where $C$ is the vector of centrality and $\bm{w}=(w_{1},...,w_{M})^{\top}$ are node weights. This weight vector can be customized based on the specific context of the research. For example, it can be defined as the ratio of liabilities to assets of the institutions or the market capitalizations of the firms. Incorporating weights allows us to assign varying degrees of importance to each node when assessing the system’s centrality. As highlighted by Ali et al., (2016), conventional centrality measures may not fully capture the systemic importance of financial firms as the size and structure of their balance sheets play a crucial role in understanding the potential risks they pose to the entire system. Larger values in this centrality measure indicate a higher level of potential spillover, particularly during times of crisis.

### v2

@@ Abstract
In this paper, we introduce an impact centrality measure to evaluate shock propagation on financial networks capturing a notion of contagion and systemic risk contributions, permitting comparisons of these risks over time. In addition, we provide a statistical validation method when the network is estimated from data, as is done in practice. This statistical test allows us to reliably assess the computed centrality values. We validate our methodology on simulated data and conduct empirical case studies using financial data. We find that our proposed centrality measure increases significantly during times of financial distress and is able to provide insights into the (market implied) risk-levels of different firms and sectors.
Keywords: Contagion Risk, Financial Stability, Network Analysis, Statistical Validation
@@ 1 Introduction
The growing complexity of financial systems requires the creation of a framework that can assess the interconnections among firms, see e.g. (Feinstein and Hałaj, , 2023, Figure 1). These connections play a crucial role in capturing systemic events and shocks that can lead to system instability. Researchers have shown growing interest in modeling financial system using network representations to gain a deeper understanding of these effects (Huang et al., , 2013; Anderson et al., , 2019; Jackson and Pernoud, , 2020; Agosto et al., , 2020). Emerging as a significant driver of instability, contagion, often described as the propagation of information or shocks across nodes, results in cascading effects that spread throughout the entire network, akin to an epidemic (Jalili and Perc, , 2017). Within this paper, we construct a network centrality measure designed to capture financial contagion. Due to the proposed construction, we are able to, e.g., statistically test our centrality measure; as far as the authors are aware, this is the first network centrality that can be analytically [^1] For some centrality measures, confidence intervals can be calculated using bootstrapping which is computationally demanding, see e.g. Antonakakis et al., (2020). tested, as opposed to other papers which test the connectivity of the network and then build the centrality measures (Billio et al., , 2012; Gabauer et al., , 2023; Basu et al., , 2024).
@@
Due to the recent financial events, it is important to view the financial system as a complex network, where firms serve as nodes and financial dependencies as links (Battiston et al., 2012a, ). There are multiple perspectives on financial networks, e.g., (i) informational (Fink et al., , 2016), (ii) interbank or default contagion (Acemoglu et al., , 2016), (iii) portfolio overlap or price-mediated contagion (Cont and Schaanning, , 2017), and (iv) cross-ownership (Fichtner et al., , 2017). The construction of financial networks involves various methods, including correlation (Pacreau et al., , 2021), partial correlation (Kenett et al., , 2015; Millington and Niranjan, , 2020), principal components analysis (Ballester et al., , 2016), and analysis of financial dependencies based on balance sheets and contracts between institutions (Eisenberg and Noe, , 2001). Studies have demonstrated that interconnectedness between institutions can be represented using weighted links, which indicate correlations between their portfolios (Diebold and Yılmaz, , 2014; Cabrales et al., , 2017) and that interpretable networks can be built on linear models (Diebold and Yilmaz, , 2009; Sudarshan Kumar and Chakrabarti, , 2022).
@@
Failures or contagion can be triggered by common factors, such as market declines, or by idiosyncratic shocks affecting individual institutions, like spread widening (Benoit et al., , 2017). The consequences of a shock can be broken down into two parts: the immediate failure of an institution triggered by a sufficiently large shock, and the spread of the shock throughout the system due to network effects. While many studies focus on the direct collapse of a firm (fundamental default) resulting from an external shock, they often overlook the aggregated impact of network-induced contagion (default by contagion) as noted by Cont et al., (2013) and Deng et al., (2021). It is important to distinguish between the first-order effect, which could be the default of a firm or the occurrence of a shock impacting the firm, and the higher-order impacts due to these stresses. As highlighted by di Iasio et al., (2013); Veraart, (2020), distress contagion can propagate through the network, causing widespread losses, even when there are no defaults within financial system. To measure these effects, it is essential to define a suitable centrality measure.
@@
Centrality measures serve as valuable tools to quantitatively assess the structural significance of nodes within a network. As stated by D’Arcangelis and Rotundo, (2016), a higher centrality score indicates that a particular property is more fitting for a given node. Different centrality measures offer distinct insights into various centrality dimensions (Hevey, , 2018). In addition, the centrality notion becomes more crucial when the networks are not symmetric (Capponi and Larsson, , 2015) which is the case in financial systems. When examining propagation within networks, Pacreau et al., (2021) highlights that the centrality of a node indicates its sensitivity to market fluctuations and reveals its connectivity within the network. Nodes with high centrality are often connected to many other nodes or are neighbors to densely interconnected subnetworks. In related work, Ghanbari et al., (2018) examines the relationship between the number of failed nodes and centrality measures. Along this line, Battiston et al., 2012a introduces DebtRank, utilizing feedback-centrality to measure the impact of a shock on a single node. Building on this concept, Puliga et al., (2014) presents Group Debtrank to quantify the centrality of nodes when small shocks affect firms and concludes that during crisis periods, the centrality measure tends to increase. Bardoscia et al., (2015) extends the framework to generalized DebtRank, with applications explored in Ferracci and Cimini, (2022) and Carro and Stupariu, (2022). Alexandre et al., (2021) applies differential DebtRank to the Brazilian interbank market, while Lv et al., (2023) introduces GuaranteeRank to evaluate default risk in intercorporate credit guarantee networks. Furthermore, there are centrality indices meaning the index is a (mainly linear) combination of different centrality measures, see e.g. Martinez-Jaramillo et al., (2014); Wang and Huang, (2021). For a literature on centrality measures refer to Bloch et al., (2023).
@@
It is crucial to select an appropriate centrality measure that aligns with the attribute under consideration (Singh et al., , 2020). In this work, we introduce a simple centrality measure that reflects the spread of a shock over the network structure based on the Leontief inverse (Leontief, , 1936, 1986). Widely used in input-output analysis in economics (Antràs et al., , 2012; Antràs and Chor, , 2013, 2018), the Leontief inverse has also found applications in network science more recently (Moran and Bouchaud, , 2019). Certain studies in the financial network field have employed the Leontief inverse to establish the network structure directly (Elliott et al., , 2014; D’Arcangelis and Rotundo, , 2016), rather than applying it to compute the centrality.
@@ 1.1 Primary Contributions
In this paper, we construct weighted directed networks, where nodes represent firms and the links depict interconnections between them. Herein we utilize the Katz-Bonacich (KB) centrality (Katz, , 1953; Bonacich, , 1972), referred to as KB, to quantify node centrality during shock scenarios, illustrating a node’s position when hit by a shock. Notably, it considers the propagation and contagion risks inherent in the network topology, and boils down the shock propagation dynamics to a closed form. However, rather than solely utilizing this analytical form of this centrality measure, we add on a statistical test layer to determine if, e.g., the determined centrality is merely an artifact of the data.
We define KB at pair-, node-, and system-levels, offering different granularities for analysis. Pair-level KB between two firms demonstrates the first firm’s position when the second firm experiences a shock. Node-level KB of a firm is the total resulting shock when it is stressed. System-level KB aggregates multiple scenarios to portray overall centrality changes under equally probable situations. We also allow the association of weights with node- and system-level centralities.
Our approach involves statistical analysis to assess the robustness of the centrality measure. We derive the asymptotic distribution of KB using a Taylor expansion and formulate hypotheses to test its significance—specifically, whether KB significantly differs from zero and whether its value varies statistically across different nodes. Since centrality measures can be quite noisy, a statistically validated result could indicate no significant difference from zero, see e.g.  (Kuzubaş et al., , 2014, Figure 14). We validate our framework using simulated data, demonstrating the theoretical distribution’s compatibility with empirical observations. In this simulated setup, the true value of KB consistently falls within the confidence intervals of the theoretical distribution at a 97.5% confidence level.
Applicability in financial contexts is a key aspect of our paper. We evaluate the performance of our framework using two financial datasets – Credit Default Swap Index (CDX) and equity tick data. In the CDX case study, KB effectively captures underlying market dynamics, outperforming both degree centrality and the leading eigenvalue. In the case of tick data, KB provides insights into the market and institutions’ states but does not suffice for standalone investment decisions. However, a blend of node- and system-level KB dynamics along with institutional background information can guide investment strategies effectively.
The structure of this paper is as follows: Section 2 outlines the approach used to construct the network and introduces the Katz-Bonacich centrality measure. With this definition, we detail the statistical framework for the KB in Section 3 and validate this framework using simulated data. We apply the KB approach to financial data scenarios in Section 4 to assess framework’s applicability to capture stress events. Finally, in Section 5, we summarize the key findings of the paper and propose potential avenues for future research.
@@ 2 Network Centrality
A network consists of a set of nodes and links. The connectivity pattern between nodes is captured by the $M\times M$ real-valued adjacency matrix $A\in\mathbb{R}^{M\times M}$, where $M$ denotes the number of nodes, and its elements $a_{ij}$, for $i,j=1,...,M$, indicate the presence or absence of connections between nodes. For our purpose, the elements $a_{ij}$ represent the connectivity weights between nodes $i$ and $j$ (Papana et al., , 2017), specifically capturing the propagation of shocks between these nodes. In this paper, we consider the nodes to be the firms and the links to be the connections.
Assumption 1 (Stationarity condition).
@@
For the remainder of this work, we assume that the leading eigenvalue of the adjacency matrix is strictly less than one.
@@
We introduce a centrality measure that as opposed to eigenvector centrality (see, e.g., Markose et al., (2012); Chan-Lau, (2018); Olmo, (2021)) but similar to DebtRank (Battiston et al., 2012b, ), is comparable over time. This measure defined as in Equation pair-level KB below, is readily interpretable in the context of financial contagion.
2.1 Katz-Bonacich Centrality Measure
Our centrality measure is derived from Katz-Bonacich centrality (KB) (Katz, , 1953; Bonacich, , 1972) defined as $KB_{\alpha}=(I-\alpha A)^{-1}\mathbf{1}$ where $\alpha>0$ is a decay parameter, $I$ is the identity matrix and $\mathbf{1}$ is a vector of ones. Notably, when $\alpha<\frac{1}{\lambda_{m}}$, where $\lambda_{m}$ denotes the leading eigenvalue of the adjacency matrix $A$, then the matrix inverse $(I-\alpha A)^{-1}$ is the Leontief inverse of $\alpha A$. Following Assumption 1, we can restrict $\alpha\in(0,1]$. If $\alpha\in(0,1)$ then this decay parameter assigns higher weight to shorter paths (direct relations). However, $\alpha=1$ provides a natural decay to shock propagation through the network. [^2] As the leading eigenvalue of $A$ is less than one, higher powers of $A$ contribute progressively less. This property ensures that indirect effects naturally decay in magnitude. For this reason, for the remainder of this work we will solely consider $\alpha=1$ so as to consider the unadulterated network and denote $KB:=KB_{1}=(I-A)^{-1}\mathbf{1}$.
@@ 2.1 Katz-Bonacich Centrality Measure
Under Assumption 1, the pair-level KB is well-defined and can be simplified via the Leontief inverse:
$\sum_{t=1}^{\infty}A^{t}=(I-A)^{-1}-I\,,$ (pair-level KB)
@@
KB is not solely influenced by the strength of the links but also by the overall structure of the network. A densely connected network with comparatively weaker links may yield a higher centrality measure than a sparsely connected network with stronger links.
To compute the weighted node-level centrality when node $i$ is shocked, we consider the weighted summation of each row of the centrality measure matrix:
$\bm{c}=[\,(I-A)^{-1}-I\,]\,\bm{w}\,,$ (node-level KB)
where $\bm{c}$ is the vector of centrality and $\bm{w}=(w_{1},...,w_{M})^{\top}$ are node weights. This weight vector can be customized based on the specific context of the research. For example, it can be defined as the ratio of liabilities to assets of the institutions or the market capitalizations of the firms. Incorporating weights allows us to assign varying degrees of importance to each node when assessing the system’s centrality. As highlighted by Ali et al., (2016), conventional centrality measures may not fully capture the systemic importance of financial firms as the size and structure of their balance sheets play a crucial role in understanding the potential risks they pose to the entire system. Larger values in this centrality measure indicate a higher level of potential spillover, particularly during times of crisis.
Finally, we define the system-level centrality as the total sum of all node-level centralities $\bm{c}_{i}$, in line with the approach outlined by Battiston et al., 2012b . This value provides a useful proxy for the total systemic risk captured in financial data.
$\sum_{i=1}^{M}\bm{c}_{i}\,.$ (system-level KB)
2.2 Financial Intuition of KB
To present the KB measure more intuitively and in shock scenarios, in Figure 1 we present a schematic network diagram along with its shock propagation dynamics, intuitively similar to Impulse Response literature in econometrics (Koop et al., , 1996; Pesaran and Shin, , 1998). Consider four time series, corresponding to four nodes denoted as $W,X,Y$ and $Z$, which are connected as illustrated in the figure on the left. The shock propagation pattern is shown in the figure on the right side, under the scenario in which node $W$ experiences shock $\epsilon$ at time $t=0$. Due to the connections with nodes $X$ and $Y$, this shock subsequently propagates and affects these nodes at time $t=1$. Following the network, the shock then cycles back to node $W$ and hits $Z$ at $t=2$. This cascading effect goes on until the shock dissipates.
[…]
In order to formulate the cumulative impact of a shock more generally, we assume $M$ time series and define an initial shock vector $\bm{\epsilon}\in\mathbb{R}^{M}$, where element $i$ represents the magnitude of the shock to node $i$. The shock propagation dynamics can be shown as:
$\begin{gathered}\bm{s}_{t}=\bm{s}_{t-1}A,\quad\bm{s}_{0}=\bm{\epsilon}\quad\Rightarrow\quad\bm{s}_{t}=\bm{\epsilon}A^{t},\end{gathered}$
where $\bm{s}_{t}$ is the state of the system at time $t$ and $A$ is the adjacency matrix. By applying the adjacency matrix $A$ to each state iteratively, we can trace the propagation of the shock over time. Specifically, $\bm{s}_{1}=\bm{s}_{0}A$, $\bm{s}_{2}=\bm{s}_{1}A=\bm{s}_{0}A^{2}$, and so on. Each term in the series $\bm{\epsilon}A^{t}$ represents the effect of the shock propagating through the network over different time intervals, accounting for higher-order interactions and contagion effects. The summation
$\sum_{t=1}^{\infty}\bm{s}_{t}=\bm{\epsilon}A+\bm{\epsilon}A^{2}+...=\bm{\epsilon}\,\underbrace{(A+A^{2}+...)}_{\text{pair-level KB}}$
captures the cumulative effect of the shock propagation and represents the centrality of the system. Each element in this matrix represents the pair-level KB, i.e., the centrality of node $j$ (corresponding to the column index) when node $i$ (corresponding to the row index) is hit by a shock. Note that this pair-level KB is independent of the initial shock $\bm{\epsilon}$.